OpenStudy (anonymous):

what is the square root of -64?

OpenStudy (anonymous):

imaginary numbers \[\sqrt{-64}\] = \[\sqrt{64}\times \sqrt{-1}\] = \[8\times i \] \[(i = \sqrt{-1})\]

OpenStudy (anonymous):

so I assume it would be D.) not a real number

OpenStudy (anonymous):

we cannot take normal roots of negative numbers, so here is how we go about it: \[\sqrt{-64} = \sqrt{-1 * 64} = \sqrt{-1} * \sqrt{64} = i * 8 = 8i\]

OpenStudy (anonymous):

Yeah, its not a real number

OpenStudy (anonymous):

thx you !